# Evolutionary algorithms meet self-supervised learning: a comprehensive survey

### v1

@@ Abstract
The number of studies that combine Evolutionary Machine Learning and self-supervised learning has been growing steadily in recent years. Evolutionary Machine Learning has been shown to help automate the design of machine learning algorithms and to lead to more reliable solutions. Self-supervised learning, on the other hand, has produced good results in learning useful features when labelled data is limited. This suggests that the combination of these two areas can help both in shaping evolutionary processes and in automating the design of deep neural networks, while also reducing the need for labelled data. Still, there are no detailed reviews that explain how Evolutionary Machine Learning and self-supervised learning can be used together. To help with this, we provide an overview of studies that bring these areas together. Based on this growing interest and the range of existing works, we suggest a new sub-area of research, which we call Evolutionary Self-Supervised Learning and introduce a taxonomy for it. Finally, we point out some of the main challenges and suggest directions for future research to help Evolutionary Self-Supervised Learning grow and mature as a field.
Index Terms:
Evolutionary Self-Supervised learning, Evolutionary Machine Learning, Self-supervised learning, Neuroevolution, Deep LearningI Introduction
Evolutionary Machine Learning (EML) [1, 2, 3, 4] is a field that has attracted the attention of researchers due to its ability to produce unexpected Machine Learning (ML) models that solve a given problem. This ability comes from the fact that EML algorithms are stochastic and, therefore, able to promote the emergence of robust solutions under noisy conditions. As the field evolves, recent developments have demonstrated competitive or superior performance compared to traditional methods, particularly in terms of generalisation, adaptability, interpretability and efficiency.
@@
The growing interest in these fields has sparked a new line of research that merges EML and SSL. This combination can bring benefits by increasing the robustness of the representations learned by SSL algorithms or improving the evolution process behind the learned solutions, especially in situations where labelled data is limited.
To the best of our knowledge, no survey focuses on works that intersect EML and SSL with such a broad view. Nevertheless, we have identified a survey that focuses specifically on the use of Evolutionary Computation (EC) in the context of Generative Adversarial Networks (GANs) [10]. Since a dedicated survey already covers this topic, we exclude it from our scope and direct readers to it for further details. In this paper, we define what we call Evolutionary Self-Supervised learning (E-SSL) as a new area of research. In total, we identified 72 papers that fall within the scope of this work. The number of these publications by year is shown in Figure 1, and it supports the idea that interest in this area is growing. To find these papers, we used search terms related to EC together with terms related to SSL, using the AND operator. Some of the EC terms include: evolutionary, evolutionary neural architecture search, genetic programming, evolutionary strategies, and cma-es. For the SSL keywords, we used terms like self-supervised, pretext task, autoencoder, contrastive, siamese networks, and unsupervised pretraining. We ran our searches on Google Scholar, IEEExplore, ACM Digital Library, and SpringerLink.
The main contributions of this survey are: (i) The definition of a new area of research called Evolutionary Self-Supervised Learning, along with a taxonomy, based on a review of recent work; (ii) A discussion of open challenges in the field and an outline of possible directions for future work within the proposed taxonomy.
The remainder of this paper is organised as follows: Section II gives background on the SSL field, while Section III provides an overview of the EML field; Section IV describes the new research area and the proposed taxonomy; Section V covers open challenges and future research directions; Finally, Section VI wraps up the paper by summarising the main contributions.
II Self-Supervised Learning (SSL)
@@
The SSL paradigm is particularly useful when the number of labelled inputs is limited, as it can be divided into two distinct stages. In the first stage, a model is trained without using any labels and without explicitly considering the final goal. Instead, pseudo-labels are used to supervise training, which forces the model to learn how to extract features for a task that is related to the actual goal but not the same. This initial step is known as the pretext task (or proxy task). The second stage, called the downstream task, uses the model trained in the pretext task to help solve the actual target task. This stage follows a supervised learning setting, where features extracted from labelled inputs are used to train the model that solves the target task. The representations learned during the pretext task can either be reused directly or fine-tuned during this second step. An overview of this two-step process in SSL is shown in Figure 2.
@@
III Evolutionary Machine Learning (EML)
@@
As for the third category, it brings ML into EC. One possible role of ML in an Evolutionary Algorithm (EA) is to improve specific parts of the evolutionary process – such as genotype-to-phenotype mapping, fitness evaluation, variation operators, parent selection, or replacement strategy.
For instance, the fitness function can be replaced by an ML model that assigns the fitness of an individual [45], use LLMs to perform “intelligent” crossover [46], or learn new genotype-phenotype mappings [47]. Additionally, EA components come with parameters, such as the probability of applying crossover or mutation operators, and the population size. Since these parameters affect the evolutionary process, ML can be used to select suitable values – either during evolution or a priori. ML can also be applied to the outputs of an EA to summarise, filter, or select results.
IV Evolutionary Self-Supervised Learning (E-SSL)
E-SSL is defined as a field that focuses on the application of EC to SSL, and vice versa. Some aspects of the EML taxonomy proposed by Banzhaf et al. [1] can be reused in this context, as the relationship between EC and SSL is similar to the one observed between EC and ML.
However, since SSL is typically divided into two stages, this opens the possibility of introducing evolution at different points in the learning process. Based on this, E-SSL works are grouped into two main categories: those that use EC to support SSL (Section IV-A), and those that apply SSL within EC algorithms (Section IV-B).
IV-A EC for SSL
The application of EC algorithms to SSL can be viewed along two distinct dimensions. The first concerns the stage of the learning process that EC targets. As noted in Section II, SSL is typically divided into two tasks – pretext and downstream – and EC can be applied to either. The second dimension relates to the specific component of the SSL algorithm being optimised.
@@ IV-A EC for SSL
IV-A1 Pretext task – Dataset
@@
A limited number of studies focus on evolving components that affect the dataset used in the pretext task. Among the works identified, two main groups can be distinguished: one where the EA optimises the pseudo-labels assigned to the data, and another where the EA targets the inputs. In the latter case, EAs can operate directly on the inputs – through generation or combinatorial optimisation – or act on a function that modifies the original data.
@@
In the first group, Li et al.[18] apply a EA to explore which combinations of pseudo-labels are most effective for representation learning in the GenNAS framework. GenNAS trains a network to learn representations by having its feature maps approximate synthetic signals, as illustrated in Figure3. A convolutional layer is added to each stage of the network, and a loss function is designed to match the output of each auxiliary layer to a synthetic signal. The authors use a Genetic Algorithm (GA) to evolve parameters that generate the synthetic signals at each stage, evaluating the quality of the learned representations on a small set of network architectures. The assumption is that the closer the network’s output is to the evolved target signal at each stage, the better the solution.
@@
Barrett et al. [48] evolve data augmentation components in the image domain. They assume that, given a set of available augmentation functions, the intensity values of these functions can be optimised during the pretext task. The authors encode both the augmentation functions and their intensity levels, and use an EA to find the optimal data augmentation configuration based on performance in the downstream task.
@@
IV-A2 Pretext task – Topology
A significant portion of E-SSL studies focus on optimising model topology using EC, as the structure of the model plays a key role in task performance. One of the most important factors when evolving topologies is the size of the search space, as it affects how an EA converges toward optimal regions. In this context, Tabak et al. [49] constrain the evolution of AEs in several ways. They use a GA with a variable-length evolutionary representation, where each gene encodes the number of neurons in a layer (restricted to dense layers only).
@@ IV-A2 Pretext task – Topology
Suganuma et al. [51] use a (1+)-ES to evolve convolutional AEs. To reduce the search space, the approach follows a symmetric design, where the decoder is a mirrored version of the encoder. The evolutionary process is restricted to convolutional layers, and both layer dimensionality and kernel size are encoded, with values selected from a predefined set. However, the EA also encodes the connectivity between layers, offering greater flexibility beyond traditional sequential architectures.
@@
Finally, Dimanov et al. [57] introduce MONCAE, a multi-objective EA that evolves convolutional AEs based on reconstruction error and compression ability. After evolution, an additional fine-tuning phase is applied by training the final population for more epochs. Individuals exceeding a predefined compression threshold are then evaluated in a downstream image classification task using three datasets: MNIST, F-MNIST, and CIFAR-10.
@@
Xue et al. [62] evolve new architectures from an existing overparameterised network that supports multiple layer options. Mutations are retained only if they result in lower validation loss, which is computed through contrastive self-supervised learning.
In a similar direction, MaskTAS [63] evolves vision transformer architectures within a supernet-constrained search space. Prior to the NAS stage, a supernet is trained using masked image modelling and self-distillation. In the first step, a teacher supernet is trained on masked image patches, following the approach of He et al.[31] (see SectionII). Then, a student supernet is trained by sampling subnets from the supernet at each batch. These subnets are updated using both the masking task and a prediction loss based on latent representations from the teacher.
During the NAS phase, an EA samples subnets from the student supernet and evolves them. Individuals are evaluated using a similarity metric that compares their learned representations to those from the teacher. The top k individuals are selected as parents, and variation is introduced through crossover and mutation. This process continues until a sufficient number of valid individuals – meeting supernet constraints – are generated. The best individual is then fine-tuned to obtain final performance results.
@@
Topologies do not necessarily need to be neural networks. Structures evolved from EAs can behave as ML models and replace deep neural networks. Within this spectrum, Rodriguez-Coayahuitl et al. [65] evolve two forest of GP structures. Two populations are maintained, one to encode the inputs and the other to decode the output. At the macro level, the algorithm behaves like a GA and each element from the GA contains a tree-based representation. The overview of their AE is depicted in figure 6.
@@
IV-A3 Pretext task – Learning
The earliest approaches in this category design EAs to evolve the weights of an AE, positioning EAs as an alternative to gradient descent and removing the requirement for differentiable neural network architectures.

### v2

# Evolutionary Machine Learning meets Self-Supervised Learning: a comprehensive survey

@@ Abstract
Research that combines Evolutionary Machine Learning and Self-Supervised Learning has been steadily increasing in recent years. Evolutionary Machine Learning has been shown to help automate the design of machine learning algorithms and to lead to more reliable solutions. Self-Supervised Learning, on the other hand, has produced good results in learning useful features when labelled data is limited. This suggests that the combination of these two areas can help both in shaping evolutionary processes and in automating the design of deep neural networks, while also reducing the need for labelled data. However, no detailed surveys exist that explain how Evolutionary Machine Learning and Self-Supervised Learning can be used together. To help with this, we provide an overview of studies that bring these areas together. Driven by the growing interest in this topic and the range of existing works, we consolidate current knowledge on approaches that combine these two fields, propose a definition for this area of research, which we call Evolutionary Self-Supervised Learning, and introduce a taxonomy for it. Finally, we point out some of the main challenges and suggest directions for future research to help Evolutionary Self-Supervised Learning grow and mature as a field.
keywords:
Evolutionary Self-Supervised Learning , Evolutionary Machine Learning , Self-Supervised Learning , Neuroevolution , Deep Learning[affiliation] organization=University of Coimbra, CISUC/LASI, DEI, city=Coimbra, country=Portugal
1 Introduction
Evolutionary Machine Learning (EML) [1, 2, 3, 4] is a field that has attracted the attention of researchers due to its ability to produce competitive, and often unexpected, Machine Learning (ML) models that solve a given problem. This ability comes from the fact that EML algorithms are stochastic and, therefore, able to promote the emergence of robust solutions under noisy conditions. Additionally, since they don’t rely on human-made solutions, they may find unexpected ones. As the field evolves, recent developments have demonstrated competitive or superior performance compared to traditional methods, particularly in terms of generalisation, adaptability, interpretability and efficiency.
@@ 1 Introduction
The growing interest in these fields has sparked a line of research that merges EML and SSL. This combination can bring benefits by increasing the robustness of the representations learned by SSL algorithms or improving the evolution process behind the learned solutions, especially in situations where labelled data is limited.
The surveys about EML that were identified [2, 3, 4] do not handle the topic of leveraging unlabelled data for representation learning despite listing a few studies that fit the subject. Similarly, surveys on SSL [10, 11] fall short in terms of detailing the efforts to automate the search or design of pretext tasks. Interestingly, conda activate neat covers the intersection of Supervised Learning with EML [12]. However, to the best of our knowledge, no survey focuses on works that intersect EML and SSL with such a broad view. Nevertheless, we have identified a survey that focuses specifically on the use of Evolutionary Computation (EC) in the context of Generative Adversarial Networks (GANs) [13]. Since a dedicated survey already covers this topic, we exclude it from our scope and direct readers to it for further details. In the scope of this paper, we focus on studies that combine EML and SSL, defining a field which encompasses all the works that fit within this intersection, naming it Evolutionary Self-Supervised Learning (E-SSL). In total, we identified 72 papers that fall within the scope of this work. The number of these publications by year is shown in Figure 1, and it supports the idea that interest in this area is growing. To find these papers, we used search terms related to EC together with terms related to SSL, using the AND operator. Some of the EC terms include: evolutionary, evolutionary neural architecture search, genetic programming, evolutionary strategies, and cma-es. For the SSL keywords, we used terms like self-supervised, pretext task, autoencoder, contrastive, siamese networks, and unsupervised pretraining. We ran our searches on Google Scholar, IEEExplore, ACM Digital Library, and SpringerLink.
The main contributions of this survey are: (i) the formalisation of Evolutionary Self-Supervised Learning (E-SSL) as a research area, together with a taxonomy derived from recent studies; and (ii) a discussion of open challenges and promising directions for future work within this taxonomy.
The remainder of this paper is organised as follows: Section 2 gives background on the SSL field, while Section 3 provides an overview of the EML field; Section 4 describes the new research area and the proposed taxonomy; Section 5 covers open challenges and future research directions; Finally, Section 6 wraps up the paper by summarising the main contributions.
2 Self-Supervised Learning (SSL)
@@
The SSL paradigm is particularly useful when the number of labelled inputs is limited, as it can be divided into two distinct stages. In the first stage, a model is trained without using any labels and without explicitly considering the final goal. Instead, pseudo-labels are used to supervise training, which forces the model to learn how to extract features for a task that is related to the actual goal but not the same. This initial step is known as the pretext task. The second stage, called the downstream task, uses the model trained in the pretext task to help solve the actual target task. This stage follows a Supervised Learning setting, where features extracted from labelled inputs are used to train the model that solves the target task. The representations learned during the pretext task can either be reused directly or fine-tuned during this second step. An overview of this two-step process in SSL is shown in Figure 2.
@@
3 Evolutionary Machine Learning (EML)
@@
As for the third category, it brings ML into EC. One possible role of ML in EC is to improve specific parts of the evolutionary process – such as genotype-to-phenotype mapping, fitness evaluation, variation operators (e.g. mutation and crossover), parent selection, or replacement strategy.
For instance, the fitness function can be replaced by an ML model that assigns the fitness of an individual [48], use LLMs to perform “intelligent” crossover [49], or learn new genotype-phenotype mappings [50]. Additionally, components from EC algorithms come with parameters, such as the probability of applying crossover or mutation operators, and the population size. Since these parameters affect the evolutionary process, ML can be used to select suitable values – either during evolution or a priori. ML can also be applied to the outputs of an EC algorithm to summarise, filter, or select results.
4 Evolutionary Self-Supervised Learning (E-SSL)
E-SSL can be seen as a subset of EML. Concretely, it concerns the application of EC to ML models that learn through SSL. The stochastic element brought by EC is a factor that can enhance the quality of the representations learned during the pretext task, as well as exploit any of their potential shortcomings. Moreover, EC is able to do the same for the downstream task.
In order to contextualise E-SSL and relate it with other fields, a visual diagram that explains how E-SSL is positioned against related areas like ML, EC and EML is depicted in Figure 3. In this figure, one can see the conceptual overlaps and distinctions among these fields.
Since we consider E-SSL to be a subset of EML, some aspects of the EML taxonomy proposed by Banzhaf et al. [1] can be reused in this context. However, since SSL is typically divided into two stages, this opens the possibility of introducing evolution at different points in the learning process. Based on this, E-SSL works are grouped into two main categories: those that use EC to support SSL (Section 4.1), and those that apply SSL within EC algorithms (Section 4.2).
4.1 EC for SSL
The application of EC algorithms to SSL can be viewed along two distinct dimensions. The first concerns the stage of the learning process that EC targets. As noted in Section 2, SSL is typically divided into two tasks – pretext and downstream – and EC can be applied to either. The second dimension relates to the specific component of the SSL algorithm being optimised.
@@ 4.1 EC for SSL
The works that fall under this section are described in Table 1. These studies are characterised in terms of components targeted by EC, and how the evolution engages with SSL. More specifically, we identify the type of SSL goals defined for the problem, whether the SSL objective type is directly customised by the evolution process, the stage at which EC is applied. Additionally, we focus on identified how the evolution process evaluated the candidate solution, as well as highlighting whether the work developed by the authors conducted test to assess representation robustness in scenarios of label scarcity.
| EC engagement with SSL | Evaluation | ||||||
| Work name | What is evolved | SSL goal type | SSL customised by EC | How | When is EC applied | Fitness metric | Label scarcity study |
| CGP-NASV2 [51] | Encoder | Rotation prediction | No | - | During pretext | Pretext accuracy | No |
| EvoDeNSS [52] | Encoder, Learning hyperparams | Feature correlation | No | - | During pretext | Downstream accuracy | Yes |
| ELo [24] | Loss function | Multiple (combined) | Yes | Linear combination of SSL losses | During pretext | Zipf law fit metric | Yes |
| Barrett et al. [53] | Data augmentation policy | Multiple (separately) | No | - | During downstream | Downstream accuracy | No |
| CSENet [54] | Loss function | Multiple (combined) | Yes | Linear combination of SSL losses | During pretext | Downstream accuracy | No |
| Li et al. [55] | Loss function, Data augmentation policy | Multiple (separately) | No | - | During pretext | Agent Score | No |
| FaUNAE [56] | Encoder | Contrastive | No | - | During pretext | Validation loss (pretext) | No |
| AutoSSL [57] | Loss function | Multiple (combined) | Yes | Linear combination of SSL losses | During pretext | Homophily based on cluster assignments | No |
| AutoNAS [58] | Encoder | Input reconstruction | No | - | During pretext | Validation loss (pretext) | No |
| Shen et al. [59] | Learning hyperparams | Rotation prediction + Color arrangement prediction (combined) | No | - | During downstream | Not reported | No |
| GenNAS [21] | Data labels | Feature maps prediction | Yes | Generation of pseudo-labels | During pretext | Validation loss (Pretext) | No |
| GenNAS with Eproxy [60] | Data labels, Learning hyperparams, Projector | Feature maps prediction | Yes | Generation of pseudo-labels | During pretext | Validation loss (Pretext) | No |
| RLNAS [22] | Encoder | Random label prediction | No | - | During pretext | Convergence based metric | No |
| StGP [61] | Downstream network | Multiple (separately) | No | - | During downstream | F1 score | No |
| David and Greental [62] | Encoder | Input reconstruction | No | - | During pretext | Train loss (Pretext) | No |
| EvoAE [63] | Encoder | Input reconstruction | No | - | During pretext | Validation loss (Pretext) | No |
| Assunção et al. [64] | Encoder, Projector | Input reconstruction | No | - | During pretext | Downstream accuracy + Net structure metric | No |
| EvoVAE [65] | Encoder, Projector | Input reconstruction | No | - | During pretext | Train loss (pretext) | Yes |
| EUDNN [66] | Encoder, Projector | Evolved | Yes | Weights evolved | During pretext | Downstream accuracy | No |
| SEvoAE [67] | Encoder | Input reconstruction | No | - | During pretext | L1 validation loss + L2 validation loss (Pretext) | No |
| EC engagement with SSL | Evaluation | ||||||
| Work name | What is evolved | SSL goal type | SSL customised by EC | How | When is EC applied | Fitness metric | Label scarcity study |
| Ludwig and Claes [68] | Encoder | Masking + Contrastive | No | - | During downstream | Word Error Rate | No |
| MaskTAS [69] | Encoder | Masking + Distillation | No | - | During pretext | Feature similarity between student and teacher | No |
| Kanwal et al. [70] | Encoder | Input reconstruction | No | - | During pretext | Validation loss (pretext) + Downstream accuracy + Net structure | No |
| ST-ITO [71] | Downstream coefficients | Audio effect and preset prediction | No | - | During downstream | Cosine similarity | No |
| Sors et al. [72] | Loss function | Contrastive | Yes | Linear combination of loss components | During pretext | Downstream mean average precision | No |
| LTR-CE [72] | Data sample pairs | Masking | No | - | Before downstream | Validation loss (downstream) | No |
| CLAMP-ViT [73] | Encoder | Contrastive | No | - | After pretext | Validation loss (pretext and downstream) | No |
| Evol-Q [74] | Encoder | Contrastive | No | - | After pretext | Validation loss (downstream) | No |
| LPQ [75] | Encoder | Contrastive | No | - | After pretext | Validation loss (pretext) + Net structure | No |
| JCLB [76] | Downstream network | Contrastive + Masking | No | - | During downstream | Downstream accuracy | No |
| AOC-VAE [77] | Encoder, Projector | Input reconstruction | No | - | During pretext | Validation loss (pretext) | Yes |
| ADWSL [78] | Learning hyperparams | Input reconstruction | No | - | During pretext | Train loss (pretext) | Yes |
| FCAE [79] | Encoder, Projector | Input reconstruction | No | - | During pretext | Train loss (pretext) | Yes |
| MONCAE [80] | Encoder, Projector | Input reconstruction | No | - | During pretext | Validation loss (pretext) | No |
| Hajewski and Oliveira [81, 82] | Encoder, Projector | Input reconstruction | No | - | During pretext | Validation loss (pretext) | No |
| Charte et al. [83, 84] | Encoder, Projector, Loss function | Input reconstruction | No | - | During pretext | Train loss (pretext) | No |
| eVAE [85] | Loss function | Input reconstruction | No | - | During pretext | Train loss (pretext) | No |
| MP2 [86] | Prompts | unknown | No | - | During downstream | unknown | No |
| SA-eInfoVAE [87] | Loss function | Input reconstruction | No | - | During pretext | Train loss (pretext) | No |
| EC engagement with SSL | Evaluation | ||||||
| Work name | What is evolved | SSL goal type | SSL customised by EC | How | When is EC applied | Fitness metric | Label scarcity study |
| Tabak et al. [88] | Encoder | Input reconstruction | No | - | During pretext | Loss (Pretext) | No |
| Preen et al. [89] | Encoder, Learning hyperparams | Input reconstruction | No | - | During pretext | Downstream accuracy | No |
| Bu et al. [90] | Learning hyperparams | Masking | No | - | During downstream | Predicted validation accuracy from surrogate model | No |
| Hajewski et al. [91, 92] | Encoder | Input reconstruction | No | - | During pretext | Validation loss (pretext) | No |
| GP autoencoder [93] | Encoder, Projector | Input reconstruction | No | - | During pretext | Loss (pretext) | No |
| GPE-AE [94] | Encoder | Input reconstruction | No | - | During pretext | Loss (pretext) | No |
| EvoMAE [95] | Encoder, Downstream network | Input reconstruction | No | - | During pretext + During downstream | Validation accuracy (Downstream) | No |
| E-CAE [96] | Encoder, Projector | Input reconstruction | No | - | During pretext | Peak Signal to Noise ratio on validation set (pretext) | No |
| SAENE [97] | Encoder, Projector | Input reconstruction | No | - | During pretext | Local Continuity Meta-Criterion metric on validation set (pretext) | No |
4.1.1 Pretext task – Dataset
@@
A limited number of studies focus on evolving components that affect the dataset used in the pretext task. Among the works identified, two main groups can be distinguished: one where the EC algorithm optimises the pseudo-labels assigned to the data, and another where evolution targets the inputs. In the latter case, the evolutionary process can operate directly on the inputs – through generation or combinatorial optimisation – or act on a function that modifies the original data.
@@
In the first group, Li et al. [21] apply a EC algorithm to explore which combinations of pseudo-labels are most effective for representation learning in the GenNAS framework. GenNAS trains a network to learn representations by having its feature maps approximate synthetic signals, as illustrated in Figure4. A convolutional layer is added to each stage of the network, and a loss function is designed to match the output of each auxiliary layer to a synthetic signal. The authors use a Genetic Algorithm (GA) to evolve parameters that generate the synthetic signals at each stage, evaluating the quality of the learned representations on a small set of network architectures. The assumption is that the closer the network’s output is to the evolved target signal at each stage, the better the solution.
@@
Barrett et al. [53] evolve data augmentation components in the image domain. They assume that, given a set of available augmentation functions, the intensity values of these functions can be optimised during the pretext task. The authors encode both the augmentation functions and their intensity levels, and use an EC algorithm to find the optimal data augmentation configuration based on performance in the downstream task.
@@
4.1.2 Pretext task – Topology
A significant portion of E-SSL studies focus on optimising model topology using EC, as the structure of the model plays a key role in task performance. One of the most important factors when evolving topologies is the size of the search space, as it affects how an EC algorithm converges toward optimal regions. In this context, Tabak et al. [88] constrain the evolution of AEs in several ways. They use a GA with a variable-length evolutionary representation, where each gene encodes the number of neurons in a layer (restricted to dense layers only).
@@ 4.1.2 Pretext task – Topology
Suganuma et al. [96] use a (1+)-ES to evolve convolutional AEs. To reduce the search space, the approach follows a symmetric design, where the decoder is a mirrored version of the encoder. The evolutionary process is restricted to convolutional layers, and both layer dimensionality and kernel size are encoded, with values selected from a predefined set. However, the evolutionary process also encodes the connectivity between layers, offering greater flexibility beyond traditional sequential architectures.
@@
Finally, Dimanov et al. [80] introduce MONCAE, a multi-objective EC algorithm that evolves convolutional AEs based on reconstruction error and compression ability. After evolution, an additional fine-tuning phase is applied by training the final population for more epochs. Individuals exceeding a predefined compression threshold are then evaluated in a downstream image classification task using three datasets: MNIST, F-MNIST, and CIFAR-10.
@@
Xue et al. [56] evolve new architectures from an existing overparameterised network that supports multiple layer options. Mutations are retained only if they result in lower validation loss, which is computed through contrastive SSL.
In a similar direction, MaskTAS [69] evolves vision transformer architectures within a supernet-constrained search space. Prior to the NAS stage, a supernet is trained using masked image modelling and self-distillation. In the first step, a teacher supernet is trained on masked image patches, following the approach of He et al. [34] (see Section2). Then, a student supernet is trained by sampling subnets from the supernet at each batch. These subnets are updated using both the masking task and a prediction loss based on latent representations from the teacher.
During the NAS phase, an EC algorithm samples subnets from the student supernet and evolves them. Individuals are evaluated using a similarity metric that compares their learned representations to those from the teacher. The top k individuals are selected as parents, and variation is introduced through crossover and mutation. This process continues until a sufficient number of valid individuals – meeting supernet constraints – are generated. The best individual is then fine-tuned to obtain final performance results.
@@
Topologies do not necessarily need to be neural networks. Structures evolved from some evolutionary process can behave as ML models and replace deep neural networks. Within this spectrum, Rodriguez-Coayahuitl et al. [93] evolve two forest of GP structures. Two populations are maintained, one to encode the inputs and the other to decode the output. At the macro level, the algorithm behaves like a GA and each element from the GA contains a tree-based representation. The overview of their AE is depicted in figure 7.
@@
4.1.3 Pretext task – Learning
The earliest approaches in this category design EC algorithms to evolve the weights of an AE, positioning them as an alternative to gradient descent and removing the requirement for differentiable neural network architectures.